# Pure Message Passing Can Estimate Common Neighbor for Link Prediction

### v3

@@ 1 Introduction
Link prediction is a cornerstone task in the field of graph machine learning, with broad-ranging implications across numerous industrial applications. From identifying potential new acquaintances on social networks (Liben-Nowell & Kleinberg, 2003) to predicting protein interactions (Szklarczyk et al., 2019), from enhancing recommendation systems (Koren et al., 2009) to completing knowledge graphs (Zhu et al., 2021), the impact of link prediction is felt across diverse domains. Recently, with the advent of Graph Neural Networks (GNNs) (Kipf & Welling, 2017) and more specifically, Message-Passing Neural Networks (MPNNs) (Gilmer et al., 2017), these models have become the primary tools for tackling link prediction tasks. Despite the resounding success of MPNNs in the realm of node and graph classification tasks (Kipf & Welling, 2017; Hamilton et al., 2018; Veličković et al., 2018; Xu et al., 2018), it is intriguing to note that their performance in link prediction does not always surpass that of simpler heuristic methods (Hu et al., 2021).
Zhang et al. (2021) highlights the limitations of GNNs/MPNNs for link prediction tasks arising from its intrinsic property of permutation invariance. Owing to this property, isomorphic nodes invariably receive identical representations. This poses a challenge when attempting to distinguish links whose endpoints are isomorphic nodes. As illustrated in Figure 1, nodes $v_{1}$ and $v_{3}$ share a Common Neighbor $v_{2}$, while nodes $v_{1}$ and $v_{5}$ do not. Ideally, due to their disparate local structures, these two links $(v_{1},v_{3})$ and $(v_{1},v_{5})$ should receive distinct predictions. However, the permutation invariance of MPNNs results in identical representations for nodes $v_{3}$ and $v_{5}$, leading to identical predictions for the two links. As Zhang et al. (2021) asserts, such node-level representation, even with the most expressive MPNNs, cannot capture structural link representation such as Common Neighbors (CN), a critical aspect of link prediction.
[…]
@@
In this work, we posit that the pure Message Passing paradigm (Gilmer et al., 2017) can indeed capture structural link representation by exploiting orthogonality within the vector space. We begin by presenting a motivating example, considering a non-attributed graph as depicted in Figure 1. In order to fulfill the Message Passing’s requirement for node vectors as input, we assign a one-hot vector to each node $v_{i}$, such that the $i$-th dimension has a value of one, with the rest set to zero. These vectors, viewed as signatures rather than mere permutation-invariant node representations, can illuminate pairwise relationships. Subsequently, we execute a single iteration of message passing as shown in Figure 1, updating each node’s vector by summing the vector of its neighbors. This process enables us to compute CN for any node pair by taking the inner product of the vectors of the two target nodes.
At its core, this naive method employs an orthonormal basis as the node signatures, thereby ensuring that the inner product of distinct nodes’ signatures is consistently zero. While this approach effectively computes CN, its scalability poses a significant challenge, given that its space complexity is quadratically proportional to the size of the graph. To overcome this, we draw inspiration from DotHash (Nunes et al., 2023) and capitalize on the premise that the family of vectors almost orthogonal to each other swells exponentially, even with just linearly scaled dimensions (Kainen & Kůrková, 1993). Instead of relying on the orthogonal basis, we can propagate these quasi-orthogonal (QO) vectors and utilize the inner product to estimate the joint structural information of any node pair.
@@ 2 Preliminaries and Related Work
Alongside the observed set of edges $E$, there exists an unobserved set of edges, which we denote as $E_{c}\subseteq V\times V\setminus E$. This unobserved set encompasses edges that are either absent from the original observation or are anticipated to materialize in the future within the graph $G$. Consequently, we can formulate the link prediction task as discerning the unobserved set of edges $E_{c}$. Heuristics link predictors include Common Neighbor (CN) (Liben-Nowell & Kleinberg, 2003), Adamic-Adar index (AA) (Adamic & Adar, 2003), and Resource Allocation (RA) (Zhou et al., 2009). CN is simply counting the cardinality of the common neighbors, while AA and RA count them weighted to reflect their relative importance as a common neighbor.
$\text{CN}(u,v)$ $=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}1;$ (1)
$\text{AA}(u,v)$ $=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}\frac{1}{\log d_{k}};$
$\text{RA}(u,v)$ $=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}\frac{1}{d_{k}}.\vspace{-4mm}$
Though heuristic link predictors are effective across various graph domains, their growing computational demands clash with the need for low latency. To mitigate this, approaches like ELPH (Chamberlain et al., 2022) and DotHash (Nunes et al., 2023) propose using estimations rather than exact calculations for these predictors. Our study, inspired by these works, seeks to further refine techniques for efficient link predictions. A detailed comparison with related works and our method is in Appendix A.
@@
The advent of graphs incorporating node attributes has caused a significant shift in research focus toward methods grounded in GNNs. Most practical GNNs follow the paradigm of the Message Passing (Gilmer et al., 2017). It can be formulated as:
${\bm{m}}_{v}^{(l)}$ $=\text{AGGREGATE}\left(\{{\bm{h}}_{v}^{(l)},{\bm{h}}_{u}^{(l)},\forall u\in\mathcal{N}_{v}\}\right),$ (2)
${\bm{h}}_{v}^{(l+1)}$ $=\text{UPDATE}\left(\{{\bm{h}}_{v}^{(l)},{\bm{m}}_{v}^{(l)}\}\right),\vspace{-10mm}$
where ${\bm{h}}_{v}^{(l)}$ represents the vector of node $v$ at layer $l$ and ${\bm{h}}_{v}^{(0)}={\bm{X}}_{v,:}$. For simplicity, we use ${\bm{h}}_{v}$ to represent the node vector at the last layer. The specific choice of the neighborhood aggregation function, $\text{AGGREGATE}(\cdot)$, and the updating function, $\text{UPDATE}(\cdot)$, dictates the instantiation of the GNN model, with different choices leading to variations of model architectures. In the context of link prediction tasks, the GAE model (Kipf & Welling, 2016) derives link representation, ${\bm{h}}{(i,j)}$, as a Hadamard product of the target node pair representations, ${\bm{h}}_{(i,j)}={\bm{h}}_{i}\odot{\bm{h}}_{j}$. Despite its seminal approach, the SEAL model (Zhang & Chen, 2018), which labels nodes based on proximity to target links and then performs message-passing for each target link, is hindered by computational expense, limiting its scalability. Efficient alternatives like ELPH (Chamberlain et al., 2022) estimate node labels, while NCNC (Wang et al., 2023) directly learns edgewise features by aggregating node representations of common neighbors.
@@ 3.1 Estimation via Mean Squared Error Regression
To explore the capacity of MPNNs in capturing the overlap information inherent in heuristic link predictors, such as CN, AA and RA, we conduct an empirical investigation, adopting the GAE framework (Kipf & Welling, 2016) with GCN (Kipf & Welling, 2017) and SAGE (Hamilton et al., 2018) as representative encoders. SEAL (Zhang & Chen, 2018), known for its proven proficiency in capturing heuristic link predictors, serves as a benchmark in our comparison. Additionally, we select a non-informative baseline estimation, simply using the mean of the heuristic link predictors on the training sets. The datasets comprise eight non-attributed graphs (more details in Section 5). Given that GNN encoders require node features for initial representation, we have to generate such features for our non-attributed graphs. We achieved this by sampling from a high-dimensional Gaussian distribution with a mean of 00 and standard deviation of $1$. Although one-hot encoding is frequently employed for feature initialization on non-attributed graphs, we choose to forgo this approach due to the associated time and space complexity.
@@ 3.2 Estimation capabilities of GNNs for link predictors
GCN: $~{}~{}\mathbb{E}{\left({\bm{h}}_{u}\cdot{\bm{h}}_{v}\right)}=\frac{C}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}\frac{1}{\hat{d}_{k}};$
SAGE: $~{}~{}\mathbb{E}{\left({\bm{h}}_{u}\cdot{\bm{h}}_{v}\right)}=\frac{C}{\sqrt{{d}_{u}{d}_{v}}}\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}1,\vspace{-4mm}$
@@
The GNN’s capability to approximate heuristic link predictors is primarily grounded in the properties of their input vectors in a linear space. When vectors are sampled from a high-dimensional linear space, they tend to be quasi-orthogonal, implying that their inner product is nearly 00 w.h.p. With message-passing, these QO vectors propagate through the graph, yielding in a linear combination of QO vectors at each node. The inner product between pairs of QO vector sets essentially echoes the norms of shared vectors while nullifying the rest. Such a trait enables GNNs to estimate CN through message-passing. A key advantage of QO vectors, especially when compared with orthonormal basis, is their computational efficiency. For a modest linear increment in space dimensions, the number of QO vectors can grow exponentially, given an acceptable margin of error (Kainen & Kůrková, 1993). An intriguing observation is that the orthogonality of QO vectors remains intact even after GNNs undergo linear transformations post message-passing, attributed to the randomized weight matrix initialization. This mirrors the dimension reduction observed in random projection (Johnson & Lindenstrauss, 1984).
@@
${\bm{h}}_{v}^{(l+1)}=\sum_{u\in\mathcal{N}_{v}}{\bm{h}}_{u}^{(l)}.\vspace{-2mm}$ (3)
Such an implementation aligns with the SAGE model that employs sum aggregation devoid of self-node propagation. This methodology also finds mention in DotHash (Nunes et al., 2023), serving as a cornerstone for our research. With this kind of message-passing design, the inner product of any node pair signatures can estimate CN impartially:
@@
$\mathbb{E}{\left({\bm{h}}_{u}\cdot{\bm{h}}_{v}\right)}$ $=\textup{CN}(u,v),$
$\mathrm{Var}{\left({\bm{h}}_{u}\cdot{\bm{h}}_{v}\right)}$ $=\frac{1}{F}\left(d_{u}d_{v}+\textup{CN}(u,v)^{2}-2\textup{CN}(u,v)\right)$
$+F\mathrm{Var}{\left({\textnormal{x}}^{2}\right)}\textup{CN}(u,v).$
@@
Both Theorem 3.1 and Theorem 3.2 underscore the significance of quasi orthogonality in input vectors, enabling message-passing to efficiently count CN. Intriguingly, in most attributed graphs, node attributes, often represented as bag-of-words (Purchase et al., 2022), exhibit inherent orthogonality. This brings forth a critical question: In the context of link prediction, do GNNs primarily approximate neighborhood overlap, sidelining the intrinsic value of node attributes? We earmark this pivotal question for in-depth empirical exploration in Appendix E, where we find that random vectors as input to GNNs can catch up with or even outperform node attributes.
@@ 4.1 QO vectors construction
Though deterministic avenues for QO vector construction are documented (Kainen, 1992; Kainen & Kurkova, 2020), our preference leans toward probabilistic techniques for their inherent simplicity. We inherit the sampling paradigm from DotHash (Nunes et al., 2023), where each node $k$ is assigned with a node signature ${\bm{h}}_{k}^{(0)}$, acquired via random sampling from the vertices of an $F$-dimensional hypercube with unit vector norms. Consequently, the sampling space for ${\bm{h}}_{k}^{(0)}$ becomes $\{-1/\sqrt{F},1/\sqrt{F}\}^{F}$.
@@
Capitalizing on the empirical observation that real-world graphs predominantly obey the power-law distribution (Barabási & Albert, 1999), we propose a strategy to control variance. Leveraging the prevalence of high-degree nodes—or hubs—we designate unique one-hot vectors for the foremost hubs. Consider the graph’s top-$b$ hubs; while other nodes draw their QO vectors from a hypercube $\{-1/\sqrt{F-b},1/\sqrt{F-b}\}^{F-b}\bigtimes\{0\}^{b}$, these hubs are assigned one-hot vectors from $\{0\}^{F-b}\bigtimes\{0,1\}^{b}$, reserving a distinct subspace of the linear space to safeguard orthogonality. Note that when new nodes are added, their QO vectors are sampled the same way as the non-hub nodes, which can ensure a tractable computation complexity.
@@
Theorem 3.1 alludes to an intriguing proposition: the estimator’s potential to encapsulate not just CN, but also RA. Essentially, RA and AA are nuanced heuristics translating to weighted enumerations of shared neighbors, based on their node degrees. In Theorem 3.2, such counts are anchored by vector norms during dot products. MPLP enhances this count methodology by rescaling node vector norms, drawing inspiration from previous works (Nunes et al., 2023; Yun et al., 2021). This rescaling is determined by the node’s representation, $\text{GNN}(u)$, and its degree $d_{u}$. The rescaled vector is formally expressed as:
@@ 4.2 Structural feature estimations
The estimator in Theorem 3.2 can effectively quantify CN. Nonetheless, solely relying on CN fails to encompass diverse topological structures embedded within the local neighborhood. To offer a richer representation, we turn to Distance Encoding (DE) (Li et al., 2020). DE acts as an adept labeling tool (Zhang et al., 2021), demarcating nodes based on their shortest-path distances relative to a target node pair. For a given pair $(u,v)$, a node $k$ belongs to a node set $\text{DE}{(p,q)}$ iff $\text{SPD}(u,k)=p$ and $\text{SPD}(v,k)=q$. Unlike its usage as node labels, we opt to enumerate these labels, producing a link feature defined by #$\text{\#}(p,q)=|\text{DE}{(p,q)}|$. Our model adopts a philosophy akin to ELPH (Chamberlain et al., 2022), albeit with a distinct node-estimation mechanism.
@@
Here, $\mathcal{N}_{v}^{s}$ pinpoints $v$’s shortest-path neighborhoods distanced by the shortest-path $s$. This method sidesteps the duplication dilemma highlighted in Theorem 3.3, ensuring that $\bm{\eta}_{v}^{(s)}$ aggregates at most one QO vector per node. Similar strategies are explored in (Abboud et al., 2022; Feng et al., 2022).
@@
The intricately designed structural features improve the expressiveness of MPLP. However, this augmented expressiveness introduces susceptibility to distribution shifts during link prediction tasks (Dong et al., 2022). Consider a scenario wherein the neighborhood of a target node pair contains a node $k$. Node $k$ resides a single hop away from one of the target nodes but requires multiple steps to connect with the other. When such a target node pair embodies a positive instance in the training data (indicative of an existing link), node $k$ can exploit both the closer target node and the link between the target nodes as a shortcut to the farther one. This dynamic ensures that for training-set positive instances, the maximum shortest-path distance from any neighboring node to the target pair is constrained to the smaller distance increased by one. This can engender a discrepancy in distributions between training and testing phases, potentially diminishing the model’s generalization capability.
To circumvent this pitfall, we adopt an approach similar to preceding works (Zhang & Chen, 2018; Yin et al., 2022; Wang et al., 2023; Jin et al., 2022). Specifically, we exclude target links from the original graph during each training batch, as shown by the dash line in Figure 3. This maneuver ensures these links are not utilized as shortcuts, thereby preserving the fidelity of link feature construction.

### v4

@@ 1 Introduction
Link prediction is a cornerstone task in the field of graph machine learning, with broad-ranging implications across numerous industrial applications. From identifying potential new acquaintances on social networks [1] to predicting protein interactions [2], from enhancing recommendation systems [3] to completing knowledge graphs [4], the impact of link prediction is felt across diverse domains. Recently, with the advent of Graph Neural Networks (GNNs) [5] and more specifically, Message-Passing Neural Networks (MPNNs) [6], these models have become the primary tools for tackling link prediction tasks. Despite the resounding success of MPNNs in the realm of node and graph classification tasks [5, 7, 8, 9], it is intriguing to note that their performance in link prediction does not always surpass that of simpler heuristic methods [10].
Zhang et al. [11] highlights the limitations of GNNs/MPNNs for link prediction tasks arising from its intrinsic property of permutation invariance. Owing to this property, isomorphic nodes invariably receive identical representations. This poses a challenge when attempting to distinguish links whose endpoints are isomorphic nodes. As illustrated in Figure 1(a), nodes $v_{1}$ and $v_{3}$ share a Common Neighbor $v_{2}$, while nodes $v_{1}$ and $v_{5}$ do not. Ideally, due to their disparate local structures, these two links $(v_{1},v_{3})$ and $(v_{1},v_{5})$ should receive distinct predictions. However, the permutation invariance of MPNNs results in identical representations for nodes $v_{3}$ and $v_{5}$, leading to identical predictions for the two links. As Zhang et al. [11] asserts, such node-level representation, even with the most expressive MPNNs, cannot capture structural link representation such as Common Neighbors (CN), a critical aspect of link prediction.
@@
In this work, we posit that the pure Message Passing paradigm [6] can indeed capture structural link representation by exploiting orthogonality within the vector space. We begin by presenting a motivating example, considering a non-attributed graph as depicted in Figure 1(a). In order to fulfill the Message Passing’s requirement for node vectors as input, we assign a one-hot vector to each node $v_{i}$, such that the $i$-th dimension has a value of one, with the rest set to zero. These vectors, viewed as signatures rather than mere permutation-invariant node representations, can illuminate pairwise relationships. Subsequently, we execute a single iteration of message passing as shown in Figure 1(b), updating each node’s vector by summing the vector of its neighbors. This process enables us to compute CN for any node pair by taking the inner product of the vectors of the two target nodes.
At its core, this naive method employs an orthonormal basis as the node signatures, thereby ensuring that the inner product of distinct nodes’ signatures is consistently zero. While this approach effectively computes CN, its scalability poses a significant challenge, given that its space complexity is quadratically proportional to the size of the graph. To overcome this, we draw inspiration from DotHash [12] and capitalize on the premise that the family of vectors almost orthogonal to each other swells exponentially, even with just linearly scaled dimensions [13]. Instead of relying on the orthogonal basis, we can propagate these quasi-orthogonal (QO) vectors and utilize the inner product to estimate the joint structural information of any node pair.
@@ 2 Preliminaries and Related Work
Alongside the observed set of edges $E$, there exists an unobserved set of edges, which we denote as $E_{c}\subseteq V\times V\setminus E$. This unobserved set encompasses edges that are either absent from the original observation or are anticipated to materialize in the future within the graph $G$. Consequently, we can formulate the link prediction task as discerning the unobserved set of edges $E_{c}$. Heuristics link predictors include Common Neighbor (CN) [1], Adamic-Adar index (AA) [14], and Resource Allocation (RA) [15]. CN is simply counting the cardinality of the common neighbors, while AA and RA count them weighted to reflect their relative importance as a common neighbor.
$\text{CN}(u,v)=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}1;~{}~{}\text{AA}(u,v)=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}\frac{1}{\log d_{k}};~{}~{}\text{RA}(u,v)=\sum_{k\in\mathcal{N}_{u}\bigcap\mathcal{N}_{v}}\frac{1}{d_{k}}.$ (1)
Though heuristic link predictors are effective across various graph domains, their growing computational demands clash with the need for low latency. To mitigate this, approaches like ELPH [16] and DotHash [12] propose using estimations rather than exact calculations for these predictors. Our study, inspired by these works, seeks to further refine techniques for efficient link predictions. A detailed comparison with related works and our method is in Appendix A.
@@
The advent of graphs incorporating node attributes has caused a significant shift in research focus toward methods grounded in GNNs. Most practical GNNs follow the paradigm of the Message Passing [6]. It can be formulated as:
${\bm{m}}_{v}^{(l)}=\text{AGGREGATE}\left(\{{\bm{h}}_{v}^{(l)},{\bm{h}}_{u}^{(l)},\forall u\in\mathcal{N}_{v}\}\right),~{}~{}{\bm{h}}_{v}^{(l+1)}=\text{UPDATE}\left(\{{\bm{h}}_{v}^{(l)},{\bm{m}}_{v}^{(l)}\}\right),$ (2)
where ${\bm{h}}_{v}^{(l)}$ represents the vector of node $v$ at layer $l$ and ${\bm{h}}_{v}^{(0)}={\bm{X}}_{v,:}$. For simplicity, we use ${\bm{h}}_{v}$ to represent the node vector at the last layer. The specific choice of the neighborhood aggregation function, $\text{AGGREGATE}(\cdot)$, and the updating function, $\text{UPDATE}(\cdot)$, dictates the instantiation of the GNN model, with different choices leading to variations of model architectures. In the context of link prediction tasks, the GAE model [17] derives link representation, ${\bm{h}}{(i,j)}$, as a Hadamard product of the target node pair representations, ${\bm{h}}_{(i,j)}={\bm{h}}_{i}\odot{\bm{h}}_{j}$. Despite its seminal approach, the SEAL model [18], which labels nodes based on proximity to target links and then performs message-passing for each target link, is hindered by computational expense, limiting its scalability. Efficient alternatives like ELPH [16] estimate node labels, while NCNC [19] directly learns edgewise features by aggregating node representations of common neighbors.
@@ 3.1 Estimation via Mean Squared Error Regression
To explore the capacity of MPNNs in capturing the overlap information inherent in heuristic link predictors, such as CN, AA and RA, we conduct an empirical investigation, adopting the GAE framework [17] with GCN [5] and SAGE [7] as representative encoders. SEAL [18], known for its proven proficiency in capturing heuristic link predictors, serves as a benchmark in our comparison. Additionally, we select a non-informative baseline estimation, simply using the mean of the heuristic link predictors on the training sets. The datasets comprise eight non-attributed graphs (more details in Section 5). Given that GNN encoders require node features for initial representation, we have to generate such features for our non-attributed graphs. We achieved this by sampling from a high-dimensional Gaussian distribution with a mean of 00 and standard deviation of $1$. Although one-hot encoding is frequently employed for feature initialization on non-attributed graphs, we choose to forgo this approach due to the associated time and space complexity.
@@ 3.2 Estimation capabilities of GNNs for link predictors
[…]
@@
The GNN’s capability to approximate heuristic link predictors is primarily grounded in the properties of their input vectors in a linear space. When vectors are sampled from a high-dimensional linear space, they tend to be quasi-orthogonal, implying that their inner product is nearly 00 w.h.p. With message-passing, these QO vectors propagate through the graph, yielding in a linear combination of QO vectors at each node. The inner product between pairs of QO vector sets essentially echoes the norms of shared vectors while nullifying the rest. Such a trait enables GNNs to estimate CN through message-passing. A key advantage of QO vectors, especially when compared with orthonormal basis, is their computational efficiency. For a modest linear increment in space dimensions, the number of QO vectors can grow exponentially, given an acceptable margin of error [13]. An intriguing observation is that the orthogonality of QO vectors remains intact even after GNNs undergo linear transformations post message-passing, attributed to the randomized weight matrix initialization. This mirrors the dimension reduction observed in random projection [20].
@@
${\bm{h}}_{v}^{(l+1)}=\sum_{u\in\mathcal{N}_{v}}{\bm{h}}_{u}^{(l)}.$ (3)
Such an implementation aligns with the SAGE model that employs sum aggregation devoid of self-node propagation. This methodology also finds mention in DotHash [12], serving as a cornerstone for our research. With this kind of message-passing design, the inner product of any node pair signatures can estimate CN impartially:
@@
$\mathbb{E}{\left({\bm{h}}_{u}\cdot{\bm{h}}_{v}\right)}$ $=\textup{CN}(u,v);$
@@
Both Theorem 3.1 and Theorem 3.2 underscore the significance of quasi orthogonality in input vectors, enabling message-passing to efficiently count CN. Intriguingly, in most attributed graphs, node attributes, often represented as bag-of-words [21], exhibit inherent orthogonality. This brings forth a critical question: In the context of link prediction, do GNNs primarily approximate neighborhood overlap, sidelining the intrinsic value of node attributes? We earmark this pivotal question for in-depth empirical exploration in Appendix E, where we find that random vectors as input to GNNs can catch up with or even outperform node attributes.
@@ 4.1 QO vectors construction
Though deterministic avenues for QO vector construction are documented [22, 23], our preference leans toward probabilistic techniques for their inherent simplicity. We inherit the sampling paradigm from DotHash [12], where each node $k$ is assigned with a node signature ${\bm{h}}_{k}^{(0)}$, acquired via random sampling from the vertices of an $F$-dimensional hypercube with unit vector norms. Consequently, the sampling space for ${\bm{h}}_{k}^{(0)}$ becomes $\{-1/\sqrt{F},1/\sqrt{F}\}^{F}$.
@@
Capitalizing on the empirical observation that real-world graphs predominantly obey the power-law distribution [24], we propose a strategy to control variance. Leveraging the prevalence of high-degree nodes—or hubs—we designate unique one-hot vectors for the foremost hubs. Consider the graph’s top-$b$ hubs; while other nodes draw their QO vectors from a hypercube $\{-1/\sqrt{F-b},1/\sqrt{F-b}\}^{F-b}\bigtimes\{0\}^{b}$, these hubs are assigned one-hot vectors from $\{0\}^{F-b}\bigtimes\{0,1\}^{b}$, reserving a distinct subspace of the linear space to safeguard orthogonality. Note that when new nodes are added, their QO vectors are sampled the same way as the non-hub nodes, which can ensure a tractable computation complexity.
@@
Theorem 3.1 alludes to an intriguing proposition: the estimator’s potential to encapsulate not just CN, but also RA. Essentially, RA and AA are nuanced heuristics translating to weighted enumerations of shared neighbors, based on their node degrees. In Theorem 3.2, such counts are anchored by vector norms during dot products. MPLP enhances this count methodology by rescaling node vector norms, drawing inspiration from previous works [12, 25]. This rescaling is determined by the node’s representation, $\text{GNN}(u)$, and its degree $d_{u}$. The rescaled vector is formally expressed as:
@@ 4.2 Structural feature estimations
The estimator in Theorem 3.2 can effectively quantify CN. Nonetheless, solely relying on CN fails to encompass diverse topological structures embedded within the local neighborhood. To offer a richer representation, we turn to Distance Encoding (DE) [26]. DE acts as an adept labeling tool [11], demarcating nodes based on their shortest-path distances relative to a target node pair. For a given pair $(u,v)$, a node $k$ belongs to a node set $\text{DE}{(p,q)}$ iff $\text{SPD}(u,k)=p$ and $\text{SPD}(v,k)=q$. Unlike its usage as node labels, we opt to enumerate these labels, producing a link feature defined by #$\text{\#}(p,q)=|\text{DE}{(p,q)}|$. Our model adopts a philosophy akin to ELPH [16], albeit with a distinct node-estimation mechanism.
@@
Here, $\mathcal{N}_{v}^{s}$ pinpoints $v$’s shortest-path neighborhoods distanced by the shortest-path $s$. This method sidesteps the duplication dilemma highlighted in Theorem 3.3, ensuring that $\bm{\eta}_{v}^{(s)}$ aggregates at most one QO vector per node. Similar strategies are explored in [27, 28].
@@
The intricately designed structural features improve the expressiveness of MPLP. However, this augmented expressiveness introduces susceptibility to distribution shifts during link prediction tasks [29]. Consider a scenario wherein the neighborhood of a target node pair contains a node $k$. Node $k$ resides a single hop away from one of the target nodes but requires multiple steps to connect with the other. When such a target node pair embodies a positive instance in the training data (indicative of an existing link), node $k$ can exploit both the closer target node and the link between the target nodes as a shortcut to the farther one. This dynamic ensures that for training-set positive instances, the maximum shortest-path distance from any neighboring node to the target pair is constrained to the smaller distance increased by one. This can engender a discrepancy in distributions between training and testing phases, potentially diminishing the model’s generalization capability.
To circumvent this pitfall, we adopt an approach similar to preceding works [18, 30, 19, 31]. Specifically, we exclude target links from the original graph during each training batch, as shown by the dash line in Figure 3. This maneuver ensures these links are not utilized as shortcuts, thereby preserving the fidelity of link feature construction.